The demand for a product for the last six years has been 15, 15, 17, 18, 20, and 19. The manager wants to predict the demand for this time series using the following simple linear trend equation: trt = 12 + 2t. What are the forecast errors for the 5th and 6th years?

A. 0, −3

B. 0, +3

C. +2, +5

D. −2, −5

E. −1, −4
Answer #1

as we know that forecast error =actual -forcasted value

forecast errors for the 5th year =20-(12+2*5)=-2

and  forecast errors for the 6th year =19-(12+2*6)=-5

option D is correct